# The $(1+3)$-dimensional ‘quantum principle of relativity’ is Einstein’s principle of relativity

### v3

@@ 2 The equivalence of boosts and ‘superboosts’
By contrast, it is not possible to determine, from Eqs. (1.1) alone, how the former are related, individually, to the new labels $y^{\prime\mu}=[c{\bf t},r^{\prime}]^{\rm T}$. There exist $4!=24$ possible identifications between the individual elements of the sets $\left\{x^{\prime\mu}\right\}$ and $\left\{y^{\prime\mu}\right\}$, and, therefore, 24 possible interpretations of DE’s proposed superboosts. Hence, in order to determine the physical content of the superboost equations (1.1), we must undo this relabeling and focus, instead, on the underlying linear transformation.
@@
But how can this be? Transformations that leave the metric invariant are symmetries of the spacetime. The symmetries of $(1+3)$-dimensional Minkowski space are well known and are described by elements of the Poincaré group [7]. This is composed of two subgroups, the group of spacetime translations and the isochronous Lorentz group, $SO^{+}(1,3)$. The Lorentz group, in turn, contains both ordinary spatial rotations and Lorentz boosts [6, 7]. Therefore, if the boosts described by Eqs. (1.1) represent symmetries of Minkowski space, then they must be elements of the Lorentz group $SO^{+}(1,3)$.
@@ 3 The $(1+1)$-dimensional limit
for any value of $1<V/c<\infty$ (or, equivalently, $01>v/c>$ 0). In this model, the local line element is not conserved under a superluminal velocity boost and the superboost defined by Eqs. (3.1) does not represent a symmetry of the spacetime. It is therefore not an element of $SO^{+}(1,3)$
@@
Nonetheless, it is conventional, in the relativity literature, to ‘absorb’ the action of the transformation matrix, $B^{\mu}{}_{\nu}(V)$, into the definition of the ‘transformed coordinates’, ${\rm d}x^{\prime\mu}=B^{\mu}{}_{\nu}(V){\rm d}x^{\nu}$. It is essential to recognise, however, that one must not transform both the basis vectors and the coordinates. This is equivalent to applying the linear map $B$ twice. Likewise, one must not apply $B$ to the basis vectors and its transpose, $B^{\rm T}$, to the components, or vice versa. This generates a passive transformation, which leaves all physical observables invariant [8]. For orthogonal matrices, this is simply equivalent to applying the identity transformation: an immediately relevant example is the transformation matrix $B^{\mu}{}_{\nu}(V\rightarrow\infty)$ (3.6), for which $B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty)=B^{\rm T}(V\rightarrow\infty)B(V\rightarrow\infty)=I$.
@@
Though we can only speculate, in the absence of explicit calculations showing their methods, it is possible that the authors’ implicit definition of the post-superboost line element, $({\rm d}s)^{2}={\rm d}{\bf x}^{\rm T}\eta{\rm d}{\bf x}\rightarrow({\rm d}s^{\prime})^{2}:=-{\rm d}{\bf x}^{\prime\rm T}\eta^{\prime}{\rm d}{\bf x}^{\prime}:={\rm d}{\bf x}^{\rm T}B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty)\eta B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty){\rm d}{\bf x}={\rm d}{\bf x}^{\rm T}\,I\,\eta\,I\,{\rm d}{\bf x}=({\rm d}s)^{2}$ (1.5) [1], misled them into believing that the metric of the spacetime could be altered without altering the local line element at any point within it. However, this is not the case, as there exists a one-to-one correspondence between the metric, $g(\mathfrak{p})=\braket{{\bf e}_{\mu}(\mathfrak{p}),{\bf e}_{\nu}(\mathfrak{p})}{\rm d}x^{\mu}(\mathfrak{p})\otimes{\rm d}x^{\nu}(\mathfrak{p})$, at a point in the spacetime manifold $\mathfrak{p}\in\mathcal{M}$, and the local line element obtained from it, $({\rm d}s)^{2}=\braket{{\bf e}_{\mu}(\mathfrak{p}),{\bf e}_{\nu}(\mathfrak{p})}{\rm d}x^{\mu}(\mathfrak{p}){\rm d}x^{\nu}(\mathfrak{p})$ [6]. Hence, in any number of spacetime dimensions, the superboost transformation $B$, enacted by the matrix $B^{\mu}{}_{\nu}({\bf V})$, transforms the canonical spacetime line element according to
@@ 4 Conclusions
    The confusion, which led the authors of [1] to believe that their $(1+3)$-dimensional theory is a generalisation of the $(1+1)$-dimensional superboost model [2], appears to have been caused by the same relabeling that obscured its equivalence to special relativity. A misunderstanding of the nature of active coordinate transformations [8], as opposed to passive coordinate transformations, also seems to have contributed.

### v4

@@ 2 The equivalence of boosts and ‘superboosts’
By contrast, it is not possible to determine, from Eqs. (1.1) alone, how the former are related, individually, to the new labels $y^{\prime\mu}=[c{\bf t}^{\prime},r^{\prime}]^{\rm T}$. There exist $4!=24$ possible identifications between the individual elements of the sets $\left\{x^{\prime\mu}\right\}$ and $\left\{y^{\prime\mu}\right\}$, and, therefore, 24 possible interpretations of DE’s proposed superboosts. Hence, in order to determine the physical content of the superboost equations (1.1), we must undo this relabeling and focus, instead, on the underlying linear transformation.
@@
But how can this be? Transformations that leave the metric invariant are symmetries of the spacetime. The symmetries of $(1+3)$-dimensional Minkowski space are well known and are described by elements of the Poincaré group [7]. This is composed of two subgroups, the group of spacetime translations and the Lorentz group, $SO(1,3)$. The Lorentz group, in turn, contains both ordinary spatial rotations and Lorentz boosts [6, 7]. Therefore, if the boosts described by Eqs. (1.1) represent symmetries of Minkowski space, then they must be elements of the Lorentz group. In particular, they must be elements of the proper isochronous Lorentz group $SO^{+}(1,3)$, as they do not include parity (P) or time-inversion (T) symmetries.
@@ 3 The $(1+1)$-dimensional limit
for any value of $1<V/c<\infty$ (or, equivalently, $01>v/c>$ 0). In this model, the local line element is not conserved under a superluminal velocity boost and the superboost defined by Eqs. (3.1) does not represent a symmetry of the spacetime. It is therefore not an element of $SO^{+}(1,3)$.
@@
Nonetheless, it is conventional, in the relativity literature, to absorb the action of the transformation matrix, $B^{\mu}{}_{\nu}(V)$, into the definition of the transformed coordinates, ${\rm d}x^{\prime\mu}=B^{\mu}{}_{\nu}(V){\rm d}x^{\nu}$. It is essential to recognise, however, that one must not transform both the basis vectors and the coordinates. This is equivalent to applying the linear map $B$ twice. Likewise, one must not apply $B$ to the basis vectors and its transpose, $B^{\rm T}$, to the components, or vice versa. This generates a passive transformation, which leaves all physical observables invariant [8]. For orthogonal matrices, this is simply equivalent to applying the identity transformation: an immediately relevant example is the transformation matrix $B^{\mu}{}_{\nu}(V\rightarrow\infty)$ (3.6), for which $B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty)=B^{\rm T}(V\rightarrow\infty)B(V\rightarrow\infty)=I$.
@@
Though we can only speculate, in the absence of explicit calculations showing their methods, it is possible that the authors’ implicit definition of the post-superboost line element, $({\rm d}s)^{2}={\rm d}{\bf x}^{\rm T}\eta{\rm d}{\bf x}\rightarrow({\rm d}s^{\prime})^{2}:=-{\rm d}{\bf x}^{\prime\rm T}\eta^{\prime}{\rm d}{\bf x}^{\prime}:={\rm d}{\bf x}^{\rm T}B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty)\eta B(V\rightarrow\infty)B^{\rm T}(V\rightarrow\infty){\rm d}{\bf x}={\rm d}{\bf x}^{\rm T}\,I\,\eta\,I\,{\rm d}{\bf x}=({\rm d}s)^{2}$ (1.5) [1], misled them into believing that the metric of the spacetime could be altered without altering the local line element at any point within it. However, this is not the case, as there exists a one-to-one correspondence between the metric, $g(\mathfrak{p})=\braket{{\bf e}_{\mu}(\mathfrak{p}),{\bf e}_{\nu}(\mathfrak{p})}{\rm d}x^{\mu}(\mathfrak{p})\otimes{\rm d}x^{\nu}(\mathfrak{p})$, at a point in the spacetime manifold $\mathfrak{p}\in\mathcal{M}$, and the local line element obtained from it, $({\rm d}s)^{2}=\braket{{\bf e}_{\mu}(\mathfrak{p}),{\bf e}_{\nu}(\mathfrak{p})}{\rm d}x^{\mu}(\mathfrak{p}){\rm d}x^{\nu}(\mathfrak{p})$ [6].
Hence, in any number of spacetime dimensions, the superboost transformation $B$, enacted by the matrix $B^{\mu}{}_{\nu}({\bf V})$, transforms the canonical spacetime line element according to
@@
Finally, it is worth noting that it is impossible to flip the metric signature of ordinary Minkowski space, in the way that DE claim, because there are no real transformations capable of achieving this. A formal proof of this statement appeared in textbooks on linear algebra as early as 1932, but was well known in the literature before that [10]. It was also well known to later generations of researchers, being referenced in the comprehensive review on superluminal transformations by Marchildon et al, in 1983 [11]. In fact, it is straightforward to see that only transformations of the form ${\rm d}x^{\mu}\rightarrow{\rm d}x^{\prime\mu}=\pm i\Lambda^{\mu}{}_{\nu}{\rm d}x^{\nu}$, where $\Lambda^{\mu}{}_{\nu}$ is a canonical Lorentz transformation, are capable of flipping the $(1+3)$-dimensional signature to the $(3+1)$-dimensional signature. To make it appear otherwise, at least superficially, one must define the signature-flipped matrix of metric components, $\eta_{\mu\nu}^{\prime}=-\eta_{\mu\nu}$ (1.4), as DE did in [1], and then replace $\eta_{\mu\nu}$ with $-\eta_{\mu\nu}^{\prime}$ in all the relevant equations (see, for example, (1.5)). Clearly, this changes nothing and is equivalent to yet another relabelling of variables, that is without physical or mathematical content.
Applying the transformation ${\rm d}x^{\mu}\rightarrow{\rm d}x^{\prime\mu}=\pm i\Lambda^{\mu}{}_{\nu}{\rm d}x^{\nu}$ to a relativistic wave equation, for example, the Klein-Gordon or Dirac equation, is equivalent to applying the transformation $\partial_{\mu}\rightarrow\partial^{\prime}_{\mu}=\pm i\partial_{\mu}$, while keeping $m$ fixed. This, in turn, is equivalent to keeping the metric fixed and mapping $m\rightarrow\pm im$, which is the standard way to define a tachyonic state [12]. In summary, the metric signature may be effectively switched, from $(1+3)$ to $(3+1)$ dimensions, using a pure-imaginary transformation that is equivalent to doing special relativity with pure-imaginary coordinates, but it cannot be flipped using DE’s formulae [1]. This is because the latter are simply the canonical Lorentz transformations, expressed in nonstandard notation.
@@ 4 Conclusions
    The confusion, which led the authors of [1] to claim that their $(1+3)$-dimensional theory is a generalisation of the $(1+1)$-dimensional superboost model [2], appears to have been caused by the same relabeling that obscured its equivalence to special relativity. A misunderstanding of the nature of active coordinate transformations [8], as opposed to passive coordinate transformations, also seems to have contributed.
Acknowledgments
This work was supported by the Grant of Scientific and Technological Projection of Guangdong Province (China), no. 2021A1515010036.